Next to a long straight wire on the amperage current I1 = 30A running through, people

install a square wire frame with current I2 = 2A. Frame and wire

straight in the same plane. The frame can rotate around a parallel axis

with the wire and go through the midpoints of two opposite sides of the frame. The shaft is rotating

one-piece conductor b = 30mm. Each frame edge has length a = 20mm. Find:

a. Force f exerts on the frame.

b. The tool is required to rotate the frame 180 ° around its axis.


1
Expert's answer
2021-01-28T12:19:05-0500

a)


"F=\frac{\mu_0 I I'a}{2\pi b}
\\F=\frac{(4\pi\cdot10^{-7})(30)(2)(0.02)}{2\pi (0.03)}=8\cdot10^{-6}N"

b)


"\tau=(8\cdot10^{-6})(0.02)=1.6\cdot10^{-7}\ Nm"